# Character Theory for Semilinear Representations

### v2

@@ 13.2. Characteristic 3
The only irreducible semilinear representation of $S_{3}$ over $L$ is the trivial representation, as $\operatorname{Irr}_{L}(H)$ consists of only the trivial representation when $L$ has characteristic $3$.
@@ 13.3. General Algorithm
It is natural to ask if one can always follow the same process used in this section to determine all $m_{K}^{L}(W)$. In general, for an orbit $\mathcal{O}$} of the action of $\Gamma$ on $\operatorname{Irr}(H,L)$ we can consider
@@ 14. Conjugacy Classes and the Number of Semilinear Representations
    $G$ is finite, $L$ is a field, with $|G|\in L^{\times}$, and
@@
This induces a natural action of $G/H\times G/H$ and hence of $\Gamma\times\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$, the subspace of functions invariant by conjugation by $H$. Because $L$ splits $H$, this has a basis given by irreducible characters $\chi\in\operatorname{Irr}(H,L)$, and with respect to this basis $\operatorname{Fun}(H/\!/H,L)$ is nothing but the permutation representation $L[\operatorname{Irr}_{L}(H)]$ of $\Gamma\times\Gamma$. Similarly, the centre $Z(L[H])$ of $L[H]$ has a basis given by conjugacy class sums $c_{h}$, which induces a permutation action of $\Gamma\times\Gamma$ on $Z(L[H])$, which is explicitly given by
@@ 17. Extension to Infinite Galois Extensions
In particular, the irreducible objects $V$ of $\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ which satisfy $L\cdot V^{N}=V$ are canonically identified with $\operatorname{Irr}_{L^{N}}^{\rtimes}(G/N)$, which is described by Theorem B.

### v3

@@ 13.2. Characteristic 3
When $L$ has characteristic $3$, $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ is no longer semisimple, and we need to understand $\operatorname{Ind}_{L}^{\rtimes}(S_{3})$. From Theorem 8.20, this is identified with $\operatorname{Ind}_{L}(H)/\operatorname{Gal}(L/K)$. The set $\operatorname{Ind}_{L}(H)$ consists of three indecomposble representations $V_{1},V_{2},V_{3}$ of dimensions $1,2,3$ respectively, and thus the action of $\operatorname{Gal}(L/K)$ on $\operatorname{Ind}_{L}(H)$ must be trivial as these dimensions are distinct. It remains to determine the semilinear Schur indices $m(V_{i})$, which we now show are all equal to $1$.
Each $V_{i}$ is defined over $K$, being given by the matrix representations where $(123)$ acts via
$(1),\qquad\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\qquad\begin{pmatrix}1&1&0\\ 0&1&1\\ 0&0&1\end{pmatrix}$
respectively. Each extends to a linear representation of $S_{3}$ over $K$, where $(12)$ acts via the matrices
$(1),\qquad\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\qquad\begin{pmatrix}2&1&0\\ 0&1&1\\ 0&0&2\end{pmatrix}$
respectively, and therefore the same matrices define extensions $\tilde{V_{i}}$ of each $V_{i}$ to $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$.
In particular, $\tilde{V_{1}},\tilde{V_{2}},\tilde{V_{3}}$ is a complete list of all indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$.
@@ 13.3. General Algorithm
It is natural to ask if one can always follow the same process used in this section to determine all $m_{K}^{L}(W)$ when $G$ is finite and $L$ has characteristic 0. In general, for an orbit $\mathcal{O}$} of the action of $\Gamma$ on $\operatorname{Irr}(H,L)$ we can consider
@@ 14. Conjugacy Classes and the Number of Semilinear Representations
    $G$ is finite, $L$ is a field, with $|H|\in L^{\times}$, and
@@
This induces a natural action of $G/H\times G/H$ and hence of $\Gamma\times\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$, the subspace of functions invariant by conjugation by $H$. Because $L$ splits $H$, this has a basis given by irreducible characters $\chi\in\operatorname{Irr}(H,L)$ [LOR, Cor. 3.21], and with respect to this basis $\operatorname{Fun}(H/\!/H,L)$ is nothing but the permutation representation $L[\operatorname{Irr}_{L}(H)]$ of $\Gamma\times\Gamma$. Similarly, the centre $Z(L[H])$ of $L[H]$ has a basis given by conjugacy class sums $c_{h}$, which induces a permutation action of $\Gamma\times\Gamma$ on $Z(L[H])$, which is explicitly given by
@@ 17. Extension to Infinite Galois Extensions
In particular, the indecomposable (resp. irreducible) objects $V$ of $\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ which satisfy $L\cdot V^{N}=V$ are canonically identified with $\operatorname{Ind}_{L^{N}}^{\rtimes}(G/N)$ (resp. $\operatorname{Irr}_{L^{N}}^{\rtimes}(G/N)$), which is described by Theorem B.